OpenStudy (anonymous):

Solve the word problem. The base of a triangle with height h and area A is given by the formula b = . What is the base of a triangle with area 17 cm2 and height 5 cm? A. 3.4 cm B. 3.8 cm C. 6.8 cm D. 11 cm

OpenStudy (anonymous):

area of a triangle=(1/2)bh plug in what you know 17=(1/2)b5 now solve for b

OpenStudy (anonymous):

How do I solve for b?

OpenStudy (anonymous):

Please help me...

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

you substitute b with your answers and see which one equals 17